OpenStudy (anonymous):

Can someone please explain how to use the half-angle formula to solve sin Pi/8?

OpenStudy (anonymous):

hi jake are yo there? do yo know the formula for half angle ?

OpenStudy (anonymous):

sin(x/2)=sqrt((1-cosx)/2)

OpenStudy (anonymous):

i think it is sin alpha/2=the square root of 1-cos alpha/2. it looks really weird having to type it out

OpenStudy (anonymous):

so alpha is just x. that makes more sense

OpenStudy (anonymous):

yes ..ok you can find x here by eqating x/2=pi/8

OpenStudy (anonymous):

i got x=.7853...

OpenStudy (anonymous):

ok use that for the right hand side

OpenStudy (anonymous):

I'm sorry I don't know what the next step is..

OpenStudy (anonymous):

ok here it is ill give you it more better

OpenStudy (anonymous):

sin((pi/4*2)=sqrt(1-cos(pi/4)/2) =sqrt((1-(sqrt2)/2))2 =sqrt((2-sqrt2))/4 =sqrt(.1464466) =.38268

OpenStudy (anonymous):

so pi/4 became your x here

OpenStudy (anonymous):

does it make sense ?

OpenStudy (anonymous):

yes that is making sense. i'm just trying to redo it on paper to see if i can get the same answer. Thank you very much

OpenStudy (anonymous):

ok wc jake

OpenStudy (anonymous):

you got it....good luck

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

k,yo have any more question or anymore prob i can help u maybe?

OpenStudy (anonymous):

are you not busy?

OpenStudy (anonymous):

not really lol

OpenStudy (anonymous):

haha well i'm in pre cal obviously and not very good at it. i have this problem: if sin theta=3/5, and tan theta<0 then in what quadrant is the point (cos theta, sin theta)?

OpenStudy (anonymous):

ok here we can put theta =x...therefore tanx=sinx/cosx so x=3/5 its in the first quadrant..but tanx is negative) so it is obvius that sin 3/5 is in the first quadrant and the cosx is in the 3rd quadrant

OpenStudy (anonymous):

cos x in the third quadrant is negative sign

OpenStudy (anonymous):

tanx <0 means tanx is a negatinve sign

OpenStudy (anonymous):

Wow thank you for your help mark